OpenStudy (anonymous):

Need Help Asap!!! Select the ordered pair from the choices below that is a solution to the following system of equations: 8x − 2 = y 5y − 12x = -3 Need Help!!!

OpenStudy (ghazi):

(5y-12x=-3) and (8x-2=y)*5 ...which will give you 40x-5y= 10 now add the equation 40x-5y=10 to 5y-12x = -3 ...you'll get x= 7/28= 1/4 hence find out y

OpenStudy (ghazi):

hope you'll get y=0

OpenStudy (anonymous):

Apreciate it for helpin out doe

OpenStudy (ghazi):

anytime..but try to give a try yourself too